OpenStudy (anonymous):

PLEASE HELP! Alg 2. Reward and medal. Need help with this I'm LOST.. 2.) A certain two-digit number has a value that is 8 less than 5 times the sum of the digits. The Tens digit is 3 less than the units digit. Find the number? (original number 10T + U) (Reversed number 10u+t) (sum (T+U)

OpenStudy (ranga):

Let the two digits be T and U The value of this number is: 10*T + U Sum of the digits is (T + U) "A certain two-digit number has a value that is 8 less than 5 times the sum of the digits." 10T + U = 5(T + U) - 8 10T + U = 5T + 5U - 8 10T - 5T = 5U - U - 8 5T = 4U - 8 ------- (1)

OpenStudy (ranga):

"The Tens digit is 3 less than the units digit." T = U - 3 ----------- (2) Solve (1) and (2)

OpenStudy (anonymous):

what does the (1) (2) ----- mean?????????

OpenStudy (ranga):

Equations one and two for easy reference later on.

OpenStudy (ranga):

Take T from equation (2) and substitute in equation (1)